# A Survey on Data Selection for LLM Instruction Tuning

### v1

@@ 3.2 Methods based on Trainable LLMs
Li et al. (2023b) present Instruction Backtranslation, a method for generating and filtering instructions. Starting with a baseline instruction-following model and a web corpus, the model generates instructions for each web document, forming a dataset. Then the model is fine-tuned with seed instructions to gain basic capabilities. And it autonomously scores each instruction, with those exceeding a set threshold forming a high-score subset for further fine-tuning. This iterative process enhances instruction generation and filtering efficiency.
@@
Li et al. (2023c) propose the Nuggets framework , which implements a dual-phase approach. Initially, it employs a variety of predefined tasks to evaluate the proficiency of a LLM across multiple scenarios, a process known as zero-shot scoring. Subsequently, each entry in the instruction dataset is utilized as a unique prompt in a one-shot manner. These prompts are presented before the predefined tasks, and the performance of LLM is reassessed, a step termed one-shot scoring. This approach harnesses the disparity between one-shot and zero-shot scores to calculate a definitive ’gold score’ for each instruction. Upon acquiring gold scores for all instructions, those constituting the highest-scoring subset are chosen as the ’gold subset’. This subset is then directly employed in the fine-tuning of the model. This methodology exploits the inherent contextual learning capabilities of extensive models.
@@
Wu et al. (2023) introduce the DIVERSEEVOL mechanism ,which represents an innovative iterative data selection strategy. It leverages large-scale models like LLaMa to generate embedding vectors for instructional data. The mechanism employs the k-center-greedy algorithmSener and Savarese (2018a) to facilitate diversity in selecting a subset of data for fine-tuning the LLaMa model. This procedure is repetitively applied, progressively enlarging the chosen subset, culminating in the creation of a high-quality instructional dataset.
@@
Chen et al. (2023c) propose the TEGIT method, offering a novel approach to generate premium instruction fine-tuning data. Particularly notable is their method of filtering instruction data. Utilizing ChatGPT, a small corpus of documentation is transformed into a format suitable for instruction data, forming a meta-dataset. This dataset then serves to train two Llama2 models - one functioning as a task generator and the other as a task discriminator. The role of generator is to devise tasks from provided texts, while the discriminator evaluates these tasks, ensuring their quality.
@@
Kung et al. (2023) present Active Instruction Tuning, which is a unique method focusing on task sensitivity selection, with the goal of enhancing large model fine-tuning using fewer tasks while improving out-of-task generalization. This technique introduces the concept of Prompt Uncertainty, which is determined by generating k perturbed instructions through the random deletion of words from an original instruction. The output of LLM probability deviation is then averaged for these k perturbed instructions. Tasks that exhibit higher Prompt Uncertainty are prioritized for instruction fine-tuning, using the extent of Prompt Uncertainty as a measure of task uncertainty.
3.3 Methods based on Powerful LLMs like ChatGPT
@@
This section introduces the method of using powerful LLMs such as GPT-4 and ChatGPT as data selectors. The approach primarily involves designing prompt templates and leveraging the capabilities of LLMs to assess the quality of instruction data.
$S_{\pi}=\{x|ChatGPT(score|prompt,x),score>\tau\}$ (9)
@@
Chen et al. (2023b) presented ALPAGASUS as an innovative data filtering methodology aimed at improving efficiency and accuracy in Instruction-Following Task (IFT) data curation. This approach utilizes a well-designed prompt applied to ChatGPT, to assess the quality of each data tuple, comprising instruction, input, and response. The methodology focuses on excluding tuples that fall below a predefined quality threshold. When this filtering process was applied to a substantial dataset, it was observed that a considerable portion of the data was compromised by quality issues. Notably, the application of the LLM-based filtering process led to the development of a model that surpassed the performance of the original model, which was trained using the unfiltered dataset with instruction-based fine-tuning.
@@
Lu et al. (2023) introduce an automated instruction tagging method (INSTAG), which employs ChatGPT to generate detailed, open-ended labels for instructions and convince the diversity and complexity of subset. The process involves annotating data with labels that reflect the semantic and intent of each instruction and normalizing them for a label-based selection approach. This approach follows a complexity-first diversification sampling strategy. It starts by sorting queries in descending order of label count and then iteratively adds queries to a subset based on the uniqueness of their labels, until the subset reaches a desired size N. The result is a curated set of queries with a higher number of labels, indicating greater complexity and diversity.
@@
To enhance the distribution and quality of datasets, Xu et al. (2023b) propose the method LIFT to reduce the redundancy of samples. It involves two stages: expanding dataset distribution and curating the diversity and quality of the dataset. Initially, ChatGPT enhances the data by generating diverse instructions and vectorizing them. A subset is then selected based on row variances. Secondly, ChatGPT scores instructions for accuracy, interpretability, clarity, difficulty, and length. The initial subset is re-selected based on these scores.
@@
Liu et al. (2023) introduce DEITA method, which integrates a multifaceted approach for selecting instruction data, focusing on complexity, quality and diversity. Complexity describes factors such as the length, difficulty, and complexity of the instruction. And quality captures the accuracy of the output. Utilizing the WizardLM technique, ChatGPT is employed to augment instructions, which are then evaluated for complexity and quality. These evaluations involve scoring instructions based on complexity, using a specially trained complexity scorer, and assessing the output quality. Each instruction in the dataset is assigned complexity scores ($c$) and quality scores ($q$), and a comprehensive score is calculated by multiplying these two metrics. The dataset is then organized according to these comprehensive scores and vectorized for further analysis.
To ensure diversity, a subset is created by adding samples that exceed a set distance threshold ($\tau$) from their nearest neighbor in the subset. This process continues until the subset reaches a predetermined size.
@@
Yu et al. (2023) present WaveCoder, a code-focused LLM enhanced through an instruction improvement technique. Its training incorporates generated data, with the data filtering phase being particularly significant. Following data generation, a LLM-based discriminator, leveraging GPT-4, evaluates instruction data against established criteria divided into subtopics. This approach enables finer control over the filtering process, effectively eliminating low-quality instruction instances.
3.4 Methods based on Small Models
@@
This section introduces methods involving the use of external small models as scorers, or the transformation of instructions into embedding vectors using small models followed by further processing. Generally, these approaches are quite comprehensive. The scoring filtering or embedding generation by Small Models often serves only as a part of the entire methodological process.
Du et al. (2023) introduce the MoDS method, focusing on instruction selection through three criteria: quality (instructional data fidelity), coverage (variety of instruction types), and necessity ( impact of instruction on LLM fine-tuning). The process unfolds in four key steps:
@@
Firstly, employing a reward model to evaluate the instruction dataset quality, selecting a subset denoted as $D_{h}$, comprising instructions that surpass a predefined quality threshold.Secondly,Utilizing the k-center-greedy algorithmSener and Savarese (2018a) to identify seed instructions, which ensures a diverse and representative sample of the instruction dataset.Thirdly,fine-tuning a pre-trained LLM with the seed instructions. Subsequently, this refined model is applied to $D_{h}$ to generate a new dataset, $D_{inference}$. This dataset is then evaluated using the reward model, which identifies instructions critical for learning of the LLM by focusing on those with lower scores. A threshold is established to select Augmented Instruction Data, which is specifically tailored to enhance performance of the model.Finally, the seed instructions with the Augmented Instruction Data to form a high-quality instruction subset that is poised to facilitate effective fine-tuning of the LLM.
@@
Chen et al. (2023a) proposes a coreset-based and task-related data selection method: firstly, obtain sentence embeddings of samples by pre-trained language models (like BERT), then select the center point by applied unsupervised clustering on these embeddings, and finally use the KCenterGreedy Sener and Savarese (2018b) algorithm to retrieve core samples from the given dataset. This method effectively reduces the required training data volume, maintaining or potentially enhancing model performance.
Table 1: Performances of different selection methods on wining rate.
@@
Selection methods Model pairs for comparison Training set Benchmark for testing (WS) Vicuna/Koala/WizardLM/Self-inst/LIMA total WS IFD Li et al. (2023a) llama-7b(5%), llama-7b(full) alpaca 1.125/0.97/1.077/1/1.1 1.04 llama-7b(10%), llama-7b(full) 1.037/1.055/1.114/1.123/1.103 1.097 llama-7b(15%), llama-7b(full) 1/1.038/1.114/1.027/1.09 1.064 llama-7b(10%), llama-7b(full) WizardLM 1.1625/1.1278/1.1147/1.0278/1.1067 1.0971 llama2-7b(5%), llama2-7b(full) alpaca 1.5875/1.4889/1.4266/1.2937/1.4733 1.4311 Random Sampling llama-7b(5%), llama-7b(full) alpaca - 0.9 InstructionGPT4 Wei et al. (2023) miniGPT4(6%), miniGPT4(full) cc_sbu_align - 1.167 Alpagsaus Chen et al. (2023b) llama-7b(9k), llama-7b(full) alpaca 1.2125/1.0222/1.0596/1.0556/- 1.0658 llama-7b(9k), llama-7b(3k) 1.1/1.183/1.082/1.17/- 1.074 llama-7b(9k), llama-7b(6k) 1.05/1.072/1.05/1.087/- 1.082 llama-13b(9k), llama-13b(full) 1.2125/1.0167/1.133/1.0198/- 1.074 MoDS Du et al. (2023) llama2-7b(2k), llama2-7b(full) alpaca 1.7125/1.5111/1.4725/1.369/1.4933 1.4786 InstructionMining Cao et al. (2023) llama-7b(2k), llama-7b(full) dolly - 1.088
@@
Table 2: Inner comparisons of LLM tuned on subsets with itself tuned on full sets.
Selection methods Training set(samples) Base model Benchmarks for testing
total WS
activeIT Kung et al. (2023) selfinstruct(2k) llama-7b 1.107
selfinstruct(full) 1.293
BBH/DROP/MMLU/Human-Eval/Avg
DQ Zhou et al. (2023b) alpaca(20%) llama-7b 32.7/26.7/39.8/9.2/27.1
alpaca(full) 32.9/26.3/41.6/10/27.7
Vicuna RS/Vicuna WTR/Koala RS/ Koala WTR
DIVERSEEVOL Wu et al. (2023) Dolly(1k) llama-7b 79.69/20/62.29/6.67
Dolly(full) 73.84/5/57.9/3.33
SelfIns(1k) 79.16/7.5/66.95/6.11
SelfIns(full) 73.03/2.5/69.5/3.89
HellaSwag/ARC/TruthfulQA/MMLU
LIFT Xu et al. (2023b) Mistral-7B 0.823/0.602/0.426/0.627
Platypus(15k random) 0.82/0.607/0.438/0.625
Platypus(15k) 0.844/0.643/0.49/0.645
RTE/CB/ANLI R1/ANLI R2/ANLI R3
coreset Chen et al. (2023a) P3(0.5) Galactica-1.3b 74.73/73.21/49.6/41.9/43.75
P3(full) 76.17/75/44/35.7/39.42
@@
Table 3: External comparisons of LLM tuned on subsets with the other LLM.
Models Training set(samples) Base model MT-bench AlpacaEval
gpt-4 8.99 95.28
gpt-3.5-turbo 7.94 91.36
alpaca-13b alpaca llama-13b 4.53 -
NUGGETSLi et al. (2023c) alpaca(7.5k) llama-7b 5.34 -
TAGLM-13b-v1.0 Lu et al. (2023) mixture(6k) llama-13b 6.44±0.04 72.8
TAGLM-13b-v2.0 Lu et al. (2023) llama2-13b 6.55±0.02 -
Instruction Length llama-13b 5.89 -
Random Sampling 5.84 -
IFDLi et al. (2023a) 5.91 -
Instruction NodeZhao et al. (2023) 5.65 -
DEITA Liu et al. (2023) mixture(10k) llama2-13b 6.79 81.09

### v2

@@ 3.2. Methods Based on Trainable LLMs
Instruction Backtranslation (Li et al., 2024b) is a method for generating and filtering instructions. Starting with a baseline instruction-following model and a web corpus, the model generates instructions for each web document, forming a dataset. Then the model is fine-tuned with seed instructions to gain basic capabilities. And it autonomously scores each instruction, with those exceeding a set threshold forming a high-score subset for further fine-tuning. This iterative process enhances instruction generation and filtering efficiency.
@@
Nuggets (Li et al., 2024a) is a framework that implements a dual-phase approach. Initially, it employs a variety of predefined tasks to evaluate the proficiency of a LLM across multiple scenarios, a process known as zero-shot scoring. Subsequently, each entry in the instruction dataset is utilized as a unique prompt in a one-shot manner. These prompts are presented before the predefined tasks, and the performance of LLM is reassessed, a step termed one-shot scoring. This approach harnesses the disparity between one-shot and zero-shot scores to calculate a definitive ’gold score’ for each instruction. Upon acquiring gold scores for all instructions, those constituting the highest-scoring subset are chosen as the ’gold subset’. This subset is then directly employed in the fine-tuning of the model. This methodology exploits the inherent contextual learning capabilities of extensive models.
@@
DIVERSEEVOL (Wu et al., 2023) is an innovative iterative data selection mechanism. It leverages large-scale models like LLaMa to generate embedding vectors for instructional data. The mechanism employs the k-center-greedy algorithm(Sener and Savarese, 2018a) to facilitate diversity in selecting a subset of data for fine-tuning the LLaMa model. This procedure is repetitively applied, progressively enlarging the chosen subset, culminating in the creation of a high-quality instructional dataset.
@@
TEGIT (Chen et al., 2023a) is a novel method for generating high-quality instruction fine-tuning data. Particularly notable is their method of filtering instruction data. Utilizing ChatGPT, a small corpus of documentation is transformed into a format suitable for instruction data, forming a meta-dataset. This dataset then serves to train two Llama2 models - one functioning as a task generator and the other as a task discriminator. The role of generator is to devise tasks from provided texts, while the discriminator evaluates these tasks, ensuring their quality.
@@
Active Instruction Tuning (Kung et al., 2023) is a method that focuses on task sensitivity selection to enhance large model fine-tuning with fewer tasks while improving out-of-task generalization. This method introduces the concept of Prompt Uncertainty, which is determined by generating k perturbed instructions through the random deletion of words from an original instruction. The output of LLM probability deviation is then averaged for these k perturbed instructions. Tasks that exhibit higher Prompt Uncertainty are prioritized for instruction fine-tuning, using the extent of Prompt Uncertainty as a measure of task uncertainty.
3.3. Methods based on Powerful LLMs like ChatGPT
@@
This section introduces the methods of using powerful LLMs such as GPT-4 and ChatGPT as data selectors. The approachs primarily involve designing prompt templates and leveraging the capabilities of LLMs to assess the quality of instruction data.
(9) $S_{\pi}=\{x|ChatGPT(score|prompt,x),score>\tau\}$
@@
ALPAGASUS (Chen et al., 2024) is an innovative data filtering methodology aimed at improving efficiency and accuracy in Instruction-Following Task (IFT) data curation. This approach utilizes a well-designed prompt applied to ChatGPT, to assess the quality of each data tuple, comprising instruction, input, and response. The methodology focuses on excluding tuples that fall below a predefined quality threshold. When this filtering process was applied to a substantial dataset, it was observed that a considerable portion of the data was compromised by quality issues. Notably, the application of the LLM-based filtering process led to the development of a model that surpassed the performance of the original model, which was trained using the unfiltered dataset with instruction-based fine-tuning.
@@
INSTAG (Lu et al., 2024) is an automated instruction tagging method that employs ChatGPT to generate detailed, open-ended labels for instructions, enhancing the diversity and complexity of the resulting subset. The process involves annotating data with labels that reflect the semantic and intent of each instruction and normalizing them for a label-based selection approach. This approach follows a complexity-first diversification sampling strategy. It starts by sorting queries in descending order of label count and then iteratively adds queries to a subset based on the uniqueness of their labels, until the subset reaches a desired size N. The result is a curated set of queries with a higher number of labels, indicating greater complexity and diversity.
@@
To enhance the distribution and quality of datasets, (Xu et al., 2023b) propose the method LIFT to reduce the redundancy of samples. It involves two stages: expanding dataset distribution and curating the diversity and quality of the dataset. Initially, ChatGPT enhances the data by generating diverse instructions and vectorizing them. A subset is then selected based on row variances. Secondly, ChatGPT scores instructions for accuracy, interpretability, clarity, difficulty, and length. The initial subset is re-selected based on these scores.
@@
DEITA (Liu et al., 2024) integrates a multifaceted approach for selecting instruction data, focusing on complexity, quality, and diversity. Complexity describes factors such as the length, difficulty, and complexity of the instruction. And quality captures the accuracy of the output. Utilizing the WizardLM technique, ChatGPT is employed to augment instructions, which are then evaluated for complexity and quality. These evaluations involve scoring instructions based on complexity, using a specially trained complexity scorer, and assessing the output quality. Each instruction in the dataset is assigned complexity scores ($c$) and quality scores ($q$), and a comprehensive score is calculated by multiplying these two metrics. The dataset is then organized according to these comprehensive scores and vectorized for further analysis. To ensure diversity, a subset is created by adding samples that exceed a set distance threshold ($\tau$) from their nearest neighbor in the subset. This process continues until the subset reaches a predetermined size.
@@
WaveCoder (Yu et al., 2023) is a code-focused LLM enhanced by an instruction improvement technique. Its training incorporates generated data, with the data filtering phase being particularly significant. Following data generation, a LLM-based discriminator, leveraging GPT-4, evaluates instruction data against established criteria divided into subtopics. This approach enables finer control over the filtering process, effectively eliminating low-quality instruction instances.
3.4. Methods based on Small Models
@@
This section introduces methods involving the use of external small models as scorers, or the transformation of instructions into embedding vectors using small models followed by further processing. Generally, these methods are quite comprehensive. The scoring filtering or embedding generation by Small Models often serves only as a part of the entire methodological process.
@@
Du et al. (2023) introduces the MoDS method, focusing on instruction selection through three criteria: quality (instructional data fidelity), coverage (variety of instruction types), and necessity ( impact of instruction on LLM fine-tuning). The process unfolds in four key steps. Firstly, employing a reward model to evaluate the instruction dataset quality, selecting a subset denoted as $D_{h}$, comprising instructions that surpass a predefined quality threshold. Secondly, utilizing the k-center-greedy algorithm (Sener and Savarese, 2018a) to identify seed instructions, which ensures a diverse and representative sample of the instruction dataset. Thirdly, fine-tuning a pre-trained LLM with the seed instructions. Subsequently, this refined model is applied to $D_{h}$ to generate a new dataset, $D_{inference}$. This dataset is then evaluated using the reward model, which identifies instructions critical for learning of the LLM by focusing on those with lower scores. A threshold is established to select Augmented Instruction Data, which is specifically tailored to enhance performance of the model.Finally, the seed instructions with the Augmented Instruction Data to form a high-quality instruction subset that is poised to facilitate effective fine-tuning of the LLM.
@@
Chen et al. (2023b) proposes a coreset-based and task-related data selection method: firstly, obtain sentence embeddings of samples by pre-trained language models (like BERT), then select the center point by applied unsupervised clustering on these embeddings, and finally use the KCenterGreedy (Sener and Savarese, 2018b) algorithm to retrieve core samples from the given dataset. This method effectively reduces the required training data volume, maintaining or potentially enhancing model performance.
3.5. Comparative Analysis of Data Selection Methods
In this subsection, we provide an in-depth comparison of the existing data selection methods across several key dimensions: model dependence, scalability, domain adaptability, interpretability, and selection effectiveness. The comparison results are shown in Table 1.
Table 1. Comparison of Instruction Data Selection Methods
Dimensions Methods in Sec.3.1 Methods in Sec.3.2 Methods in Sec.3.3 Methods in Sec.3.4
Model Dependence Model-agnostic and heuristic Requires supervised model training Depends on external black-box APIs (e.g., GPT-4) Uses small pre- trained models (e.g., BERT, TinyLM)
Scalability High Low Low Moderate
Domain Adaptability Low High Medium Low to Medium
Interpretability High Low Low Moderate
Selection Effectiveness Variable High High Medium
The first dimension, model dependence, captures how much each method relies on external or trainable models. Indicator-based approaches are generally model-agnostic, relying on heuristic or statistical features derived directly from the data, without needing to query or train large models. In contrast, trainable LLM-based methods require supervised training of a dedicated selector or reward model, often built on large-scale LLMs, making them heavily model-dependent. Methods that use powerful LLMs like GPT-4 operate in a black-box manner, scoring or filtering instructions via external APIs, which introduces reliance on proprietary infrastructure. Small model approaches strike a balance by leveraging small-scale pretrained models (e.g., BERT or TinyLM), reducing resource demands while maintaining some model-based reasoning.
In terms of scalability, indicator-based methods have a clear advantage. They are generally fast, parallelizable, and well-suited for large-scale filtering tasks, as they do not require model inference. Trainable methods, by contrast, face serious scalability challenges due to the high computational cost of model training and inference. Black-box LLM-based methods suffer from scalability bottlenecks due to API latency, usage limits, and costs, making them impractical for processing massive datasets unless sufficient resources are available. Small model methods offer a middle ground—they are not as scalable as pure heuristics but can be applied to larger datasets with reasonable efficiency.
Domain adaptability varies notably among these methods. Trainable methods exhibit the highest adaptability, as their learning process can directly align with new domains through supervised fine-tuning on relevant data. Powerful LLMs also adapt reasonably well in few-shot or zero-shot scenarios, depending on the prompt design and instruction diversity. Indicator-based methods, however, tend to be less adaptive since their features are usually handcrafted and not easily transferable across domains. Small models may offer limited adaptability depending on their pretraining corpus and whether further domain-specific fine-tuning is conducted.
When it comes to interpretability, heuristic indicator methods clearly lead, offering intuitive, transparent criteria for data selection (e.g., length, diversity, perplexity). This is important for tasks requiring human oversight or explainable decision-making. In contrast, both trainable and black-box LLM-based methods operate as opaque systems, making it difficult to trace why certain instances are selected. Small models offer moderate interpretability—they can be analyzed via model outputs and saliency techniques but are not as transparent as heuristics.
Finally, the selection effectiveness dimension compares how well each method identifies high-quality instruction data. Both trainable and powerful LLM-based methods tend to produce strong results, as they explicitly optimize or evaluate instruction usefulness using powerful language models. Small models offer moderate performance—better than pure heuristics but generally not on par with full-scale LLMs. Indicator-based methods exhibit variable performance, with effectiveness largely dependent on the specific metrics used and the task domain.
In summary, no single method dominates across all dimensions. Practitioners must balance trade-offs among efficiency, generalizability, and interpretability when choosing a data selection strategy, or consider hybrid designs that combine the strengths of multiple approaches.
4. Evaluation methods and Result Analysis
To measure the effectiveness of different instruction data selection methods, several evaluation metrics are proposed and can be divided into three categories: winning rate, inner comparison, and external comparison. In this section, we first introduce each distinct evaluation and then compare different data selection methods across multiple benchmarks under these evaluations to analyze which methods achieve better performance.
4.1. Wining Rate
To evaluate the effectiveness of the dataset selection methods, (Chen et al., 2024) define the winning rate of the LLM fine-tuned with less samples (i.e., LLM-sub) compared to the LLM fine-tuned with full samples (i.e., base LLM) and is calculated as:
(10) Win $\text{Win Rate}=\frac{Num(win)-Num(lose)}{Num(all)}+1,$
where Num(win) represents the number of wining cases, Num(lose) represents the number of losing cases, Num(all) represents the number of all cases in the testing benchmarks, and 1 indicates equal performance, values above 1 indicate improved performance, and values below 1 suggest degraded performance.
LLM-sub represents the LLM fine-tuned on the subset that filtered from the training set by the selection methods while the base LLM often involves 2 types: i) fine-tuned on the full training set and ii) fine-tuned on the same-scale subset filtered by the regular selections (e.g. random sampling and instruction length). The outputs of LLM-sub and the base LLM are rated by the judge on a scale from 1 to 10, and usually GPT4 is employed as the judge. To address the positional bias of the judge, Li et al. (2024c) send the outputs of these two LLMs to the judge twice with different orders. According to Li et al. (2024c), the wining cases refer to LLM-sub outperforming the base LLM in both times or winning in one and tying in the other. The losing cases refer to LLM-sub lagging behind the base LLM in both times or tying in one and losing in the other. The wining rates of different selection methods on the testing benchmarks are summarized in Table 2.
Table 2. Performances of different selection methods on wining rate.
@@ 4.1. Wining Rate
Selection methods Model pairs for comparison Training set Benchmark for testing (WS) Vicuna/Koala/WizardLM/Self-inst/LIMA total WS IFD (Li et al., 2024c) Llama-7B(5%), Llama-7B(full) alpaca 1.125/0.97/1.077/1/1.1 1.04 Llama-7B(10%), Llama-7B(full) 1.037/1.055/1.114/1.123/1.103 1.097 Llama-7B(15%), Llama-7B(full) 1/1.038/1.114/1.027/1.09 1.064 Llama-7B(10%), Llama-7B(full) WizardLM 1.1625/1.1278/1.1147/1.0278/1.1067 1.0971 Llama2-7B(5%), Llama2-7B(full) alpaca 1.5875/1.4889/1.4266/1.2937/1.4733 1.4311 Random Sampling Llama-7B(5%), Llama-7B(full) alpaca - 0.9 InstructionGPT4 (Wei et al., 2023) miniGPT4(6%), miniGPT4(full) cc_sbu_align - 1.167 Alpagsaus (Chen et al., 2024) Llama-7B(9k), Llama-7B(full) alpaca 1.2125/1.0222/1.0596/1.0556/- 1.0658 Llama-7B(9k), Llama-7B(3k) 1.1/1.183/1.082/1.17/- 1.074 Llama-7B(9k), Llama-7B(6k) 1.05/1.072/1.05/1.087/- 1.082 Llama-13B(9k), Llama-13B(full) 1.2125/1.0167/1.133/1.0198/- 1.074 MoDS (Du et al., 2023) Llama2-7B(2k), Llama2-7B(full) alpaca 1.7125/1.5111/1.4725/1.369/1.4933 1.4786 InstructionMining (Cao et al., 2023) Llama-7B(2k), Llama-7B(full) dolly - 1.088
@@
Table 3. Inner comparisons of LLM tuned on subsets with itself tuned on full sets.
Selection methods Training set(samples) Base model Benchmarks for testing total WS activeIT (Kung et al., 2023) selfinstruct(2k) Llama-7B 1.107 selfinstruct(full) 1.293 BBH/DROP/MMLU/Human-Eval/Avg DQ (Zhou et al., 2023b) alpaca(20%) Llama-7B 32.7/26.7/39.8/9.2/27.1 alpaca(full) 32.9/26.3/41.6/10/27.7 Vicuna RS/Vicuna WTR/Koala RS/ Koala WTR DIVERSEEVOL (Wu et al., 2023) Dolly(1k) Llama-7B 79.69/20/62.29/6.67 Dolly(full) 73.84/5/57.9/3.33 SelfIns(1k) 79.16/7.5/66.95/6.11 SelfIns(full) 73.03/2.5/69.5/3.89 HellaSwag/ARC/TruthfulQA/MMLU LIFT (Xu et al., 2023b) Mistral-7B 0.823/0.602/0.426/0.627 Platypus(15k random) 0.82/0.607/0.438/0.625 Platypus(15k) 0.844/0.643/0.49/0.645 RTE/CB/ANLI R1/ANLI R2/ANLI R3 coreset (Chen et al., 2023b) P3(0.5) Galactica-1.3B 74.73/73.21/49.6/41.9/43.75 P3(full) 76.17/75/44/35.7/39.42
@@
4.2. Inner Comparison
Inner comparison reveals the trade-off between dataset size and quality. Although models fine-tuned on the full training set generally perform well, carefully selected subset methods can achieve comparable or even better performance with fewer samples, demonstrating improved efficiency. However, the performance gains vary across different methods and tasks, highlighting the importance of adjusting selection strategies based on the characteristics of downstream tasks and the diversity of instructions.
Table 4. External comparisons of LLM tuned on subsets with the other LLM.
Methods Training set(samples) Base model MT-bench AlpacaEval gpt-4 8.99 95.28 gpt-3.5-turbo 7.94 91.36 alpaca-13B alpaca Llama-13B 4.53 - NUGGETS(Li et al., 2024a) alpaca(7.5k) Llama-7B 5.34 - TAGLM-13B-v1.0 (Lu et al., 2024) mixture(6k) Llama-13B 6.44±0.04 72.8 TAGLM-13B-v2.0 (Lu et al., 2024) Llama2-13B 6.55±0.02 - Instruction Length Llama-13B 5.89 - Random Sampling 5.84 - IFD(Li et al., 2024c) 5.91 - Instruction Node(Zhao et al., 2023) 5.65 - DEITA (Liu et al., 2024) mixture(10k) Llama2-13B 6.79 81.09